# Generalizing Visual Question Answering from Synthetic to Human-Written Questions via a Chain of QA with a Large Language Model

### v1

@@ Abstract
Visual question answering (VQA) is a task where an image is given, and a series of questions are asked about the image. To build an efficient VQA algorithm, a large amount of QA data is required which is very expensive. Generating synthetic QA pairs based on templates is a practical way to obtain data. However, VQA models trained on those data do not perform well on complex, human-written questions. To address this issue, we propose a new method called chain of QA for human-written questions (CoQAH). CoQAH utilizes a sequence of QA interactions between a large language model and a VQA model trained on synthetic data to reason and derive logical answers for human-written questions. We tested the effectiveness of CoQAH on two types of human-written VQA datasets for 3D-rendered and chest X-ray images and found that it achieved state-of-the-art accuracy in both types of data. Notably, CoQAH outperformed general vision-language models, VQA models, and medical foundation models with no finetuning.
@@ 1 Introduction
Visual question answering (VQA) aims to build an automated algorithm to answer a series of questions regarding a given image. This task has a wide range of potential applications, such as interpreting medical images Lin et al. (2023) and supporting visually impaired people Bigham et al. (2010).
Recently, general vision language models (VLMs) Alayrac et al. (2022)OpenAI (2023)Liu et al. (2023), trained using a massive amount of image and text data, have shown promising results in solving VQA tasks. However, their performance is limited when tested in VQA tasks of specific domains (e.g., VLM model trained on large amounts of medical images and captions on the web vs. VQA model specialized for chest X-rays).
@@
This challenge highlights the need to train or finetune VLM models using VQA data in a particular domain. However, acquiring such data is challenging, requiring experts to design VQA datasets carefully without contradiction and logical errors. For instance, to create a VQA dataset of chest X-rays, radiologists must write QA pairs consistent with the given radiographs.
To address this challenge, a template-based approach, which automatically synthesizes QA pairs based on pre-defined templates, has been adopted. This approach produced high-quality VQA datasets with minimum errors (e.g., CLEVR Johnson et al. (2017a) and MIMIC-Diff-VQA Hu et al. (2023)). Based on these datasets, some studies Hu et al. (2023)Hudson and Manning (2018)Perez et al. (2018)Kamath et al. (2021) reported very high accuracy on internal testing (e.g., 99.7% in CLEVR Kamath et al. (2021)). However, they also found that those VQA models failed to answer human-written questions that deviated from the templates, such as questions requiring complex reasoning (e.g., what shape is the large yellow object? vs. what shape do the most object have in common? in Figure 1b) and, even paraphrased questions (e.g., where is the lesion? vs. which area is the lesion present in? in Figure 1c).
In this study, we propose a novel method called “chain of QA for human-written questions” (CoQAH) which can correctly answer complex, human-written questions beyond a fixed set of synthetic, template-based questions without the need for finetuning. In the core part of CoQAH, a large language model (LLM) sequentially asks template-based questions to a VQA model trained with synthetic data. This step-by-step QA process enables the LLM to collect valuable and accurate information about an image and reach a reasonable conclusion for a given question.
To check the effectiveness of CoQAH, we utilized two different types of template-based VQA datasets (CLEVR Johnson et al. (2017a) for 3D-rendered images and MIMIC-Diff-VQA Hu et al. (2023) for chest X-rays) for AI training. We then used multiple human-written VQA datasets (CLEVR-Human Johnson et al. (2017b), VQA-RAD Lau et al. (2018), and SLAKE Liu et al. (2021)) for AI testing. As a result, CoQAH achieved state-of-the-art performance on all the human-written VQA datasets, surpassing general VLMs and other VQA models.
@@ 2.1 General Vision Language Models
A general VLM is an AI model that can understand images and texts, along with their relationships, in various tasks. One of the earliest approaches to creating a general VLM was to combine separate vision and text encoders, aligning features from both encoders Radford et al. (2021)Yu et al. (2022). Another approach utilized a pre-trained LLM, which can understand the meaning and context of texts. Tsimpoukelli et al. (2021) proposed training and combining a vision encoder with a frozen LLM. Similarly, Alayrac et al. (2022) proposed finetuning the connecting layers between an LLM and a vision encoder. More recently, Liu et al. (2023) proposed training VLMs using the instructions of various tasks to make a general-purpose assistant. These general VLMs can be applied to various vision-language tasks, but they often lack expertise in specific domains. CoQAH addresses this issue by employing a specialized VQA model (e.g., chest X-ray VQA) that interacts with an LLM.
@@ 2.2 Template-Based VQA
Creating VQA datasets manually is a time-consuming and expensive process. To overcome this challenge, researchers Johnson et al. (2017a)Hu et al. (2023) have attempted to synthesize QA pairs using pre-defined templates. For instance, one study Hu et al. (2023) introduced a VQA dataset for chest X-rays by extracting abnormal findings from radiologist’s reports. Similarly, another study Johnson et al. (2017a) created 3D-rendered images of various objects and generated complex template-based questions related to the visual features of those images. Several other studies Hudson and Manning (2018)Perez et al. (2018)Kamath et al. (2021)Hu et al. (2023) have used such datasets to evaluate and compare the visual understanding capabilities of AI models.
@@ 2.3 LLM-Aided VQA
Several studies Yang et al. (2022)Tiong et al. (2022)Guo et al. (2023)Zhou et al. (2023)You et al. (2023) have used an LLM to solve VQA problems. This is done by feeding the relevant information to the LLM, which then uses its high contextual understanding and reasoning ability to answer the question. For example, some researchers Yang et al. (2022)Tiong et al. (2022)Guo et al. (2023) have proposed feeding an image description from a captioning model to an LLM, while other Zhou et al. (2023) utilized VQA and captioning models together to give visual clues to an LLM. Recently, some researchers You et al. (2023) have suggested using interactions between an LLM and a VQA model to answer complex visual questions. CoQAH uses a template-based VQA model that is specialized to questions in specific domains, which makes it different from other methods above.
@@ 3.1 CoQAH
Suppose a user question and an image from an human-written VQA dataset ($q_{h}\in Q_{h},i_{h}\in I_{h};Q_{h}$ for human-written questions and $I_{h}$ for images) are given; first, the $LLM$ is prompted with a task instruction ($\rho$; see Section 3.2 for detail) that directs the model to generate a template-based question that conforms to the question templates used to train the VQA model ($q_{t}^{1}=LLM(\rho);$ $q_{t}^{1}\in Q_{t}$ where $Q_{t}$ includes all possible template-based questions; see Section 3.2). Next, $VQA_{t}$ answers the question ($a_{t}^{1}=VQA_{t}(q_{t}^{1},i_{h})$). After that, all the previous dialogue, including the task instruction, template-based questions, and answers from $VQA_{t}$ $(\rho+q_{t}^{1}+a_{t}^{1})$, is again fed into $LLM$, generating a next template-based question ($q_{t}^{2}=LLM(\rho,q_{t}^{1},a_{t}^{1})$). This QA process is repeated until $LLM$ has sufficient information to reach the final answer for the user question or it reaches the maximum number of questions to be queried (twenty questions for CLEVR-Human and five questions for VQA-RAD and SLAKE; see Section 4.4). Note that $LLM$ cannot access the image, whereas $VQA_{t}$ perceive it.
@@ 3.2 Task Instruction
The task instruction $\rho$ is a detailed description of how $LLM$ should respond to and interact with $VQA_{t}$ (Figure 2a). The instruction starts with your goal is to answer this user question based on an image: followed by the user question. Then, we included a statement to declare that $LLM$ cannot see a given image (but you can’t access the image, and I can access the image.). After that, we asked $LLM$ to give a question that can only be generated based on a template format (you can ask me questions with these forms; see Section 3.3 for details). In the last part, we added some sentences to ask $LLM$ to conclude the answer for the user question whenever it is ready (if you can answer the user question, stop asking and give me an answer.).
@@ 3.3 Template-Based Questions from LLM
Figure 2b describes how we instructed $LLM$ to generate a template-based question for the CLEVR dataset Johnson et al. (2017a). We first let $LLM$ know the structure of the question, which is composed of several different entities ($\langle question\rangle=\langle type\rangle+\langle object\rangle+\langle relation\rangle+\langle object\rangle$ in Figure 2b). Subsequently, we asked $LLM$ to choose one of the options available for each entity (e.g., small or large or $\langle Empty\rangle$ for $\langle Size\rangle$ entity). The template format for MIMIC-Diff-VQA is shown in Supplementary Figure 1. The detailed description of the task instructions for both CLEVR and MIMIC-Diff-VQA is displayed in Supplementary Figure 2 and 3.
@@ 3.4 Existence and Uniqueness Handler
To address this problem, we introduced an existence and uniqueness handler (EUH), which continually checks the violation of both conditions on a template-based question generated from $LLM$. As described in Algorithm 1, we first extracted the object entity from the question (i.e., $\langle object\rangle$ or $\langle object\rangle+\langle relation\rangle+\langle object\rangle$) and checked the existence of the object entities by querying to $VQA_{t}$ (is there a …? in Algorithm 1). When the existence of the object was confirmed, we continued to check the uniqueness by asking $VQA_{t}$ to count the number of objects (how many …? in Algorithm 1). Using EUH, we prevented $LLM$ from directly providing a wrong answer without considering logical contradiction.
@@
Input: Question from LLM $Q_{L}$, VQA model $VQA$
@@
1:  $Extract\_Entity$: Extract (Object + Relation + Object) part from the question (Type + Object + Relation + Object)
2:  $E\leftarrow Extract\_Entity(Q_{L})$
3:  $Q_{E}\leftarrow$ “Is there a” $+$ $E$ $+$ “?”
4:  $A_{E}\leftarrow VQA(Q_{E})$
5:  if $A_{E}$ is “No” then
6:     $A_{L}\leftarrow$ “There is no” $+$ $E$
7:  else if $A_{E}$ is “Yes” then
8:     $Q_{U}\leftarrow$ “How many” $+$ $E$ $+$ “are there?”
9:     $A_{U}\leftarrow VQA(Q_{U})$
10:     if $A_{U}$ is not “1” then
11:        $A_{L}\leftarrow$ “There are” $+$ $A_{U}$ $+$ $E$
12:     else if $A_{U}$ is “1” then
13:        $A_{L}\leftarrow VQA(Q_{L})$
@@ 4.1 Datasets
For datasets, we collected two types of images (and question-answer pairs), including 3D-rendered images and chest X-ray images, to show the effectiveness of CoQAH in different image types. We first trained VQA models using template-based datasets (CLEVR Johnson et al. (2017a) for 3D-rendered and MIMIC-Diff-VQA Hu et al. (2023) for chest X-ray) and externally validated the model’s performance on the human-written datasets (CLEVR-Human Johnson et al. (2017b) for 3D-rendered; VQA-RAD Lau et al. (2018) and SLAKE Liu et al. (2021) for chest X-ray).
CLEVR.
Johnson et al. (2017a) is a synthetic VQA dataset (70,000 image and QA pairs for training, 15,000 for validation, and 15,000 for testing). 3D objects of different sizes, colors, materials, and shapes were rendered for each image. Then, questions were generated based on a template-based algorithm, and the answers for those questions were automatically selected among a set of candidates (e.g., ”3”, ”blue”, ”cylinder”, etc.).
CLEVR-Human.
Johnson et al. (2017b) is a variant of the CLEVR dataset that replaced questions and answers in some images of CLEVR using human questioners (17,817 for training; 7,202 for validation). The questioners created new human-written questions without any format restriction, while the answers for the questions were chosen among the same candidates in CLEVR. Currently, a test set is not available. Therefore, we reported AI performance using the validation set.
MIMIC-Diff-VQA.
Hu et al. (2023) is a chest X-ray VQA dataset. QA pairs were generated using an automatic template algorithm based on radiologists’ reports (700,703 QA pairs from 164,324 reports and images; 80% for training, 10% for validation, and 10% for testing).
VQA-RAD and SLAKE.
Lau et al. (2018) and Liu et al. (2021) are medical image VQA datasets where clinical trainees (for VQA-RAD) or physicians (for SLAKE) wrote human-written questions and answers manually for each image. We used only the chest X-rays in those datasets (107 images for VQA-RAD and 179 images for SLAKE) for AI evaluation. In both datasets, answers are categorized as either closed-form or open-form: the answers of the closed-form questions were required to be one between two or three options (e.g., yes or no; 511 questions for VQA-RAD and 663 questions for SLAKE), whereas the answers of the open-form questions did not have any format (283 questions for VQA-RAD and 1,459 questions for SLAKE).
@@ 4.2 Metrics
[…]
4.3 Benchmarks
4.3.1 Benchmarks for CLEVR-Human
To benchmark the CLEVR-Human dataset, we employed three types of AI models, including general VLMs, VQA models trained with CLEVR (i.e., template-based VQA), and finetuned with CLEVR-Human (i.e., finetuned VQA). For the general VLMs, we utilized LLaVA Liu et al. (2023) (version 1.5; 7B and 13B parameters; source code and model weights available on GitHub; see task instruction in Supplementary Figure 4) and GPT-4-Vision OpenAI (2023) (version 1106-preview; API used; see task instruction in Supplementary Figure 5), testing them on the CLEVR-Human validation dataset. For the template-based VQA models, we benchmarked FiLM Perez et al. (2018), MAC Hudson and Manning (2018), MDETR Kamath et al. (2021), and our CoQAH. In CoQAH, we combined MDETR trained with CLEVR as a VQA model and GPT-4 as an LLM, measuring accuracy on the CLEVR-Human validation data. Finally, we compared the reported scores from the previous studies for the accuracy of the finetuned VQA models (FiLM, MAC, and MDETR) Perez et al. (2018)Hudson and Manning (2018)Kamath et al. (2021).
4.3.2 Benchmarks for VQA-RAD and SLAKE
[…]
4.3.3 Post-Processing of Answers
4.4 Other Analyses
Since the CoQAH method utilizes an LLM to understand the context and finally answer the user question, we can ask the rationale of the final answer to the LLM (i.e., interpreting the AI model’s behavior), simply querying ’why?’. For a few complex questions, we asked the LLM to explain the reasons to check whether it had logically driven their conclusions.
In addition, to validate the effect of the maximum number of questions to be queried in CoQAH, we measured the accuracy of CoQAH on the CLEVR-Human validation dataset (5% used), changing that number from 0 to 30.
As an ablation study, we excluded EUH in CoQAH and measured the accuracy based on the CLEVR-Human validation data (5% used). To check the importance of the step-by-step QA process in CoQAH, we enforced the LLM to ask all the questions simultaneously (ask me up to 20 questions all at once), measuring the accuracy. Also, we let the LLM generate any form of questions during the chain of QA process by ablating the format description in the task instruction, also calculating accuracy.
Finally, we performed an error analysis by investigating the dialogues where CoQAH failed to answer correctly (5% CLEVR-Human validation used) and categorized them according to the reasons (e.g., incorrect answers from VQA model).
5 Results
@@
LLaVA (7B) Liu et al. (2023) No No   $*{}^{*}$
LLaVA (13B) Liu et al. (2023) No No   $*{}^{*}$
GPT-4-vision OpenAI (2023) No No   $*{}^{*}$
@@
FiLM Perez et al. (2018) Yes No 56.6
MAC Hudson and Manning (2018) Yes No 57.4
MDETR Kamath et al. (2021) Yes No 59.9 ($*{}^{*}$)
CoQAH Yes No   $*{}^{*}$
@@
FiLM Perez et al. (2018) Yes Yes 75.9
MAC Hudson and Manning (2018) Yes Yes 81.5
MDETR Kamath et al. (2021) Yes Yes 81.7
Table 1: Comparison of accuracy between the general VLMs, template-based VQA models, and finetuned VQA models. CoQAH achieved the highest accuracy among all the template-based models, although it did not utilize any data from CLEVR-Human for training. * indicates accuracy measured using the CLEVR-Human validation data.
@@
BiomedGPT Zhang et al. (2023a) 43.2 0.150 - -
Med-Flamingo Moor et al. (2023) 46.4 0.184 33.2 0.140
MedVInT-TD Zhang et al. (2023b) 57.3 0.274 46.5 0.396
@@
[…]

### v3

@@ Abstract
Visual question answering (VQA) is a task where an image is given, and a series of questions are asked about the image. To build an efficient VQA algorithm, a large amount of QA data is required which is very expensive. Generating synthetic QA pairs based on templates is a practical way to obtain data. However, VQA models trained on those data do not perform well on complex, human-written questions. To address this issue, we propose a new method called chain of QA for human-written questions (CoQAH). CoQAH utilizes a sequence of QA interactions between a large language model and a VQA model trained on synthetic data to reason and derive logical answers for human-written questions. We tested the effectiveness of CoQAH on two types of human-written VQA datasets for 3D-rendered and chest X-ray images and found that it achieved state-of-the-art accuracy in both types of data. Notably, CoQAH outperformed general vision-language models, VQA models, and medical foundation models with no finetuning. The source code for CoQAH is available at https://github.com/tae2hee/CoQAH.
[…]
734
@@ 1 Introduction
Visual question answering (VQA) aims to build an automated algorithm to answer a series of questions regarding a given image. This task has a wide range of potential applications, such as interpreting medical images [12] and supporting visually impaired people [3].
Recently, general vision language models (VLMs), trained using a massive amount of image and text data, have shown promising results in solving VQA tasks [2, 1, 14]. However, their performance is limited when tested in VQA tasks of specific domains (e.g., VLM model trained on large amounts of medical images and captions on the web vs. VQA model specialized for chest X-rays) [11].
@@
This challenge highlights the need to train or finetune VLM models using VQA data in a particular domain. However, acquiring such data is challenging, requiring experts to design VQA datasets carefully without contradiction and logical errors. For instance, to create a VQA dataset of chest X-rays, radiologists must write QA pairs consistent with the given radiographs [10, 13].
To address this challenge, a template-based approach, which automatically synthesizes QA pairs based on pre-defined templates, has been adopted. This approach produced high-quality VQA datasets with minimum errors (e.g., CLEVR [7] and MIMIC-Diff-VQA [5]). Based on these datasets, some studies reported very high accuracy on internal testing (e.g., 99.7% in CLEVR) [5, 6, 17, 9]. However, they also found that those VQA models failed to answer human-written questions that deviated from the templates, such as questions requiring complex reasoning (e.g., what shape is the large yellow object? vs. what shape do the most objects have in common? in Figure 1b) and, even paraphrased questions (e.g., where is the lesion? vs. which area is the lesion present in? in Figure 1c).
In this study, we propose a novel method called chain of QA for human-written questions (CoQAH) which can correctly answer complex, human-written questions beyond a fixed set of synthetic, template-based questions without the need for finetuning. In the core part of CoQAH, a large language model (LLM) sequentially asks template-based questions to a VQA model trained with synthetic data. This step-by-step QA process enables the LLM to collect valuable and accurate information about an image and reach a reasonable conclusion for a given question.
To check the effectiveness of CoQAH, we utilized two different types of template-based VQA datasets (CLEVR for 3D-rendered images and MIMIC-Diff-VQA for chest X-rays) for AI training. We then used multiple human-written VQA datasets (CLEVR-Human [8], VQA-RAD [10], and SLAKE [13]) for AI testing. As a result, CoQAH achieved state-of-the-art performance on all the human-written VQA datasets, surpassing general VLMs and other VQA models.
@@ 2.1 General Vision Language Models
A general VLM is an AI model that can understand images and texts, along with their relationships, in various tasks. One of the earliest approaches to creating a general VLM was to combine separate vision and text encoders, aligning features from both encoders [18, 24]. Another approach utilized a pre-trained LLM, which can understand the meaning and context of texts. Tsimpoukelli et al. [20] proposed training and combining a vision encoder with a frozen LLM. Similarly, Alayrac et al. [2] proposed finetuning the connecting layers between an LLM and a vision encoder. More recently, Liu et al. [14] proposed training VLMs using the instructions of various tasks to make a general-purpose assistant. These general VLMs can be applied to various vision-language tasks, but they often lack expertise in specific domains. CoQAH addresses this issue by employing a specialized VQA model (e.g., chest X-ray VQA) that interacts with an LLM.
@@ 2.2 Template-Based VQA
Creating VQA datasets manually is a time-consuming and expensive process. To overcome this challenge, some researchers have attempted to synthesize QA pairs using pre-defined templates [5, 7]. For instance, one study introduced a VQA dataset for chest X-rays by extracting abnormal findings from radiologist’s reports [5]. Similarly, another study created 3D-rendered images of various objects and generated complex template-based questions related to the visual features of those images [7]. Several other studies have used such datasets to train and evaluate the visual understanding capabilities of AI models [6, 17, 9, 5].
@@ 2.3 LLM-Aided VQA
Several studies have used an LLM to solve VQA problems [22, 19, 4, 27, 23]. This is done by feeding the relevant information to the LLM, which then uses its high contextual understanding and reasoning ability to answer the question. For example, some researchers [22, 19, 4] have proposed feeding an image description from a captioning model to an LLM, while others have utilized VQA and captioning models together to give visual clues to an LLM [27]. Recently, some researchers have suggested using iterative interactions between an LLM and a VQA model to answer complex visual questions [23]. CoQAH uses a template-based VQA model that is specialized to questions in specific domains, which makes it different from other methods above.
@@ 3.1 CoQAH
Suppose a user question and an image from an human-written VQA dataset ($q_{h}\in Q_{h},i_{h}\in I_{h};Q_{h}$ for human-written questions and $I_{h}$ for images) are given; first, the $LLM$ is prompted with a task instruction ($\rho$; see Section 3.2 for detail) that directs the model to generate a template-based question that conforms to the question templates used to train the VQA model ($q_{t}^{1}=LLM(\rho);$ $q_{t}^{1}\in Q_{t}$ where $Q_{t}$ includes all possible template-based questions; see Section 3.2). Next, $VQA_{t}$ answers the question ($a_{t}^{1}=VQA_{t}(q_{t}^{1},i_{h})$). After that, all the previous dialogue, including the task instruction, template-based questions, and answers from $VQA_{t}$ $(\rho+q_{t}^{1}+a_{t}^{1})$, is again fed into $LLM$, generating a next template-based question ($q_{t}^{2}=LLM(\rho,q_{t}^{1},a_{t}^{1})$). This QA process is repeated until $LLM$ has sufficient information to reach the final answer for the user question or it reaches the maximum number of questions to be queried (twenty questions for CLEVR-Human and five questions for VQA-RAD and SLAKE; see Section 5.4). Note that $LLM$ cannot access the image, whereas $VQA_{t}$ does.
@@ 3.2 Task Instruction
The task instruction $\rho$ is a detailed description of how $LLM$ should respond to and interact with $VQA_{t}$ (Figure 2a). The instruction starts with Your task is to answer the following user question based on an image: followed by the user question. Then, we included a statement to declare that $LLM$ cannot see a given image (However, you cannot view the image, while I can.). After that, we asked $LLM$ to give a question that can only be generated based on a template format (You may ask me questions using the following format; see Section 3.3 for details). In the last part, we added some sentences to ask $LLM$ to query the next question step-by-step after having the response from $VQA_{t}$ (Ask me the next question after I have responded to you.) and conclude the answer for the user question whenever it is ready (Once you are able to answer the user question, stop asking further questions and provide me with the answer.).
@@ 3.3 Template-Based Questions from LLM
Figure 2b describes how we instructed $LLM$ to generate a template-based question for the CLEVR dataset [7]. We first let $LLM$ know the structure of the question, which is composed of several different entities ($\langle question\rangle=\langle type\rangle+\langle object\rangle+\langle relation\rangle+\langle object\rangle$ in Figure 2b). Subsequently, we asked $LLM$ to choose one of the options available for each entity (e.g., small or large or $\langle Empty\rangle$ for $\langle Size\rangle$ entity). The template format for MIMIC-Diff-VQA is shown in Supplementary Figure 1. The detailed description of the task instructions for both CLEVR and MIMIC-Diff-VQA is displayed in Supplementary Figure 2 and 3.
@@ 3.4 Existence and Uniqueness Handler
To address this problem, we introduced an existence and uniqueness handler (EUH), which continually checks the violation of both conditions on a template-based question generated from $LLM$. As described in Algorithm 1, we first extracted the object entity ($\langle object\rangle+\langle relation\rangle+\langle object\rangle$) from the question ($\langle type\rangle+\langle object\rangle+\langle relation\rangle+\langle object\rangle$). For example, the object entity ("small object" + "right of" + "shiny red object") was extracted from the question (What color is the small object right of the shiny red object?). Then, we checked the existence of the object entity by querying to $VQA_{t}$ (e.g., Is there a small object right of the shiny red object?). When the existence of the object was confirmed, we continued to check the uniqueness by asking $VQA_{t}$ to count the number of objects (e.g., How many small objects right of the shiny red object are there?). Using EUH, we prevented $LLM$ from directly providing a wrong answer without considering logical contradiction.
@@
Input: Question from LLM $Q_{L}$; Input image $I$; VQA model $VQA(\cdot)$; A function that extracts $\langle\textit{Object}\rangle+\langle\textit{Relation}\rangle+\langle\textit{Object}\rangle$ part from the question $\langle\textit{Type}\rangle+\langle\textit{Object}\rangle+\langle\textit{Relation}\rangle+\langle\textit{Object}\rangle$ $Extract\_Entity(\cdot)$
@@
1:  $E\leftarrow Extract\_Entity(Q_{L})$
2:  $Q_{E}\leftarrow$ “Is there a" $+$ $E$ $+$ “?"
3:  $A_{E}\leftarrow VQA(Q_{E},I)$
4:  if $A_{E}$ is “No" then
5:     $A_{L}\leftarrow$ “There is no" $+$ $E$
6:  else if $A_{E}$ is “Yes" then
7:     $Q_{U}\leftarrow$ “How many" $+$ $E$ $+$ “are there?"
8:     $A_{U}\leftarrow VQA(Q_{U},I)$
9:     if $A_{U}$ is not “1" then
10:        $A_{L}\leftarrow$ “There are" $+$ $A_{U}$ $+$ $E$
11:     else if $A_{U}$ is “1" then
12:        $A_{L}\leftarrow VQA(Q_{L},I)$
13:     end if
@@ 4.1 Datasets
For datasets, we collected two types of images (and question-answer pairs), including 3D-rendered images and chest X-ray images, to show the effectiveness of CoQAH in different image types. We first trained VQA models using template-based datasets (CLEVR [7] for 3D-rendered and MIMIC-Diff-VQA [5] for chest X-ray) and externally validated the model’s performance on the human-written datasets (CLEVR-Human [8] for 3D-rendered; VQA-RAD [10] and SLAKE [13] for chest X-ray).
CLEVR
is a synthetic VQA dataset (70,000 image and QA pairs for training, 15,000 for validation, and 15,000 for testing). 3D objects of different sizes, colors, materials, and shapes were rendered for each image. Then, questions were generated based on a template-based algorithm, and the answers for those questions were automatically selected among a set of candidates (e.g., 3, blue, cylinder, etc.).
CLEVR-Human
is a variant of the CLEVR dataset that replaced questions and answers in some images of CLEVR using human questioners (17,817 for training; 7,202 for validation). The questioners created new human-written questions without any format restriction, while the answers for the questions were chosen among the same candidates in CLEVR. Currently, a test set is not available. Therefore, we reported AI performance using the validation set.
MIMIC-Diff-VQA
is a chest X-ray VQA dataset. QA pairs were generated using an automatic template algorithm based on radiologists’ reports (700,703 QA pairs from 164,324 reports and images; 80% for training, 10% for validation, and 10% for testing).
VQA-RAD and SLAKE
are medical image VQA datasets where clinical trainees (for VQA-RAD) or physicians (for SLAKE) wrote human-written questions and answers manually for each image. We used only the chest X-rays in those datasets (107 images for VQA-RAD and 179 images for SLAKE) for AI evaluation. In both datasets, answers are categorized as either closed-form or open-form: the answers of the closed-form questions were required to be one between two or three options (e.g., yes or no; 511 questions for VQA-RAD and 663 questions for SLAKE), whereas the answers of the open-form questions did not have any formatting restrictions (283 questions for VQA-RAD and 1,459 questions for SLAKE).
@@ 4.2 Metrics
We measured exact-match accuracy for the human-written questions that select answers from predefined candidates (all questions of CLEVR-Human and closed-form questions of VQA-RAD and SLAKE). For the open-form questions of VQA-RAD and SLAKE, we calculated LAVEGPT-4 [15] that uses an LLM (GPT-4 in this paper) for the evaluation of answers.
Table 1: Comparison of accuracy between the general VLMs, template-based VQA models, and finetuned VQA models. CoQAH achieved the highest accuracy with large gaps compared to all the other general VLMs and template-based VQA models, although it did not utilize any data from CLEVR-Human for training. * indicates accuracy measured using the CLEVR-Human validation data.
@@
LLaVA (7B) [14] No No   43.8
LLaVA (13B) [14] No No   47.6
GPT-4-vision [1] No No   60.1
@@
FiLM [17] Yes No 56.6
MAC [6] Yes No 57.4
MDETR [9] Yes No 59.9 (59.5)
CoQAH Yes No   74.3
@@
FiLM [17] Yes Yes 75.9
MAC [6] Yes Yes 81.5
MDETR [9] Yes Yes 81.7
Table 2: Comparison of the performance between the medical foundation and template-based VQA models on VQA-RAD and SLAKE. CoQAH reported the highest accuracy and LAVEGPT-4 in both open- and closed-form questions. There were huge gaps between the LAVEGPT-4 of the OFA-MIMIC and CoQAH, meaning that the chain of QA process in CoQAH indeed improved the correctness of the answers significantly.
@@
BiomedGPT [25] 43.2 0.150 - -
Med-Flamingo [16] 46.4 0.184 33.2 0.140
MedVInT-TD [26] 57.3 0.274 46.5 0.396
@@
4.3 Benchmarks
4.3.1 Benchmarks for CLEVR-Human
To benchmark the CLEVR-Human dataset, we employed three types of AI models, including general VLMs, VQA models trained with CLEVR (i.e., template-based VQA), and finetuned with CLEVR-Human (i.e., finetuned VQA). For the general VLMs, we utilized LLaVA [14] (version 1.5; 7B and 13B parameters; source code and model weights available on GitHub; see task instruction in Supplementary Figure 4) and GPT-4-Vision [1] (version 1106-preview; API used; see task instruction in Supplementary Figure 5), testing them on the CLEVR-Human validation dataset. For the template-based VQA models, we benchmarked FiLM [17], MAC [6], MDETR [9], and our CoQAH. In CoQAH, we combined MDETR trained with CLEVR as a VQA model and GPT-4 as an LLM, measuring accuracy on the CLEVR-Human validation data. Finally, we compared the reported scores from the previous studies for the accuracy of the finetuned VQA models (FiLM, MAC, and MDETR).
4.3.2 Benchmarks for VQA-RAD and SLAKE
To benchmark the VQA-RAD and SLAKE datasets, we employed three general VLMs specialized for the medical domain (i.e., medical foundation models): BiomedGPT [25] (no task instruction needed), Med-Flamingo [16] (see task instruction in Supplementary Figure 6), and MedVInT-TD [26] (see task instruction in Supplementary Figure 7). As template-based VQA models, we benchmarked OFA [21] trained by ourselves (i.e., OFA-MIMIC; using a model weight of OFAbase with 182M parameters; AdamW optimizer with learning rate = 1e-4, $\beta_{1}$ = 0.9, and $\beta_{2}$ = 0.999; batch size = 16) because, so far, no VQA model trained with MIMIC-Diff-VQA is available to report performance on VQA-RAD and SLAKE. In CoQAH, we combined the OFA-MIMIC and GPT-4. Note that, different from the benchmarks on CLEVR-Human, we could not finetune the OFA-MIMIC using VQA-RAD or SLAKE because the number of chest X-rays is very limited (107 X-rays for VQA-RAD and 179 X-rays for SLAKE).
4.3.3 Post-Processing of Answers
5 Results